OpenStudy (anonymous):

What is the slope of the equation 5x + 4y = 8 ? Answer: a. 1/3 b. -5/4 c. 5/8 d. -4/5 e. 8/5

OpenStudy (anonymous):

e

OpenStudy (anonymous):

how is dat?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I apologize, I was looking at another question. The answer is b. 5x + 4y = 8 Subtract 8 5x + 4y -8 = 0 Subtract 4y 5x - 8 = -4y Divide by -4 -5/4x - 2 = y Slope is m in y = mx + b Slope= -5/4

OpenStudy (anonymous):

r yu sure